# Challenges and recommendations for Electronic Health Records data extraction and preparation for dynamic prediction modelling in hospitalized patients - a practical guide

### v1

@@ 4.4 Data cleaning
Timestamps when data are available in the system for real-time prediction differ from the clinically relevant timestamp The clinically relevant timestamp reflects the time that best corresponds to the patient’s actual physiological state, and can differ significantly from the data recording timestamp. For laboratory (and microbiology) results (Sauer et al., 2022; Tomašev et al., 2021), the relevant timestamps in the clinical workflow are: order time (when the clinician orders a lab test), planned collection time, collection execution time (when the nurse collects the sample), preliminary result availability time and final (approved) result availability time. Typically, the sample collection time and final result time are extracted from the system. For example, the time when a positive microbiology result is available in the laboratory system is often more than one day after the blood collection by the nurse; the latter is the clinically relevant timestamp because the result reflects the patient state at blood collection time (one day earlier than the result availability). Using the wrong timestamp in prediction will create temporal leaks, but its impact on the individual predictions will differ depending on how features are aggregated in time windows and how often the predictions are renewed (e.g.: predictions renewed hourly, every 24 hours or tigger based - when new data become available in the system) Most of the times (and especially for lab results), the relevant timestamps can be extracted from the system. Whenever not possible, the extraction documentation should mention the meaning of the extracted timestamp and its usual recording procedure and possible delays in recording. The registration time in the system (result time for lab results) should be used for features included in the prediction model. If the outcome is calculated retrospectively based on extracted lab results (e.g.: AKI, sepsis, CLABSI), the clinically relevant timestamp (sample collection time for lab results) should be used for outcome calculation. Timeliness - currency; Documentation Currency
The extracted timestamp does not reflect the time when data are recorded in the system. The precise timestamp when items are recorded (or even modified) in the system might not be extracted, or might not be stored in the EHR database for extraction. Typical timestamps differences occur for nurse and clinician observations and coded procedures. For example, nurses might record observations in the patient chart after they finish a shift and they might record an approximation of the time they observed the patient, e.g: at 10:13 the nurse inserts a wound care observation with the time of observation: 8:20 (an estimation of the time the nurse visited the patient). Coded procedures (CPT codes), such as catheter insertion, are typically recorded at or after the time of the procedure’s completion and the timestamp assigned to the procedure is the execution time or a close approximation of it. Often, CPT codes are documented retrospectively, even after patient discharge, when coding is reviewed for billing, based on patient’s documentation throughout the hospitalisation. There can be variation in the recording process, even within the same procedure code: sometimes procedures codes are recorded at the time of execution and other times after patient discharge. The discrepancy between the data available at model development and data available at prediction time in the clinical implementation will typically be larger for coded procedures compared to the nurse tasks. The extraction documentation should mention the meaning of the extracted timestamp and its usual recording procedure and possible delays in recording. These are though difficult to document exhaustively, as the nurse flow might differ in different wards and coding procedures might vary widely, situations in which understanding of the flow will require communication with the clinical experts. Timeliness - currency; Documentation Currency
Exact-hour timestamps often do not reflect the time when data are recorded in the system If a large proportion of timestamps in a date/time field appear as exact hours (e.g., 08:00:00 or 23:00:00), it may indicate that the extracted timestamps deviate from the time of recording in the system. For example, nurses might set the execution time of the task in the software to the current time at the moment of execution and/or execute tasks but record them in the EHR software later, leaving the task execution time to the planned time for the task, which is a generic time or the nurse shift. Depending on the EHR software and the hospital flow, it could also mean that the task was planned, but not executed (when task execution status is not available in the extraction). The data extraction documentation of the extracted timestamp columns should mention the exact meaning of the extracted timestamps: execution, planning or registration time, or a combination thereof. Although the procedures for data recording might differ from ward to ward, whenever possible, the hospital flow should be documented and understood. When the exact-hour timestamps do not represent the meaningful timestamps, but the clinical items recorded are of predictive value, encoding the timestamps at the end of the relevant time unit (e.g.: end of nurse shift) would help avoiding temporal leaks Timeliness - currency; Documentation Currency
@@
No date or timestamps available Some items are not extracted with a timestamp or have date but no time. For example, ICD codes are linked to admission and usually not timestamped; pharmacy medication orders have a date associated, but no timestamp. A conservative approach is to encode the timestamp at the end of the relevant time unit (e.g.: ICD codes at the end of the admission (Tomašev et al., 2021); medication orders end of day) or a less conservative approach by assigning a typical time of registration (e.g.: a very high percentage of medications are ordered before noon, assign 12:00 PM) Timeliness - currency Currency
@@
Temporal leaks due to extraction errors Temporal leaks can also occur because of errors during the ETL process, for example introduced by incorrect time zone handling, incorrect joins or time based filtering clauses. These are typically difficult to spot at data preparation time. If the temporal leaks have a profound effect on the prediction performance, features that are not expected to have very high importance might appear as most important (e.g.: history of bacteraemia being the most important feature in predicting bacteraemia during current admission might point to incorrectly leaking the current admission in the extraction of the historical feature). Timestamps of some measurements might exceed the admission/discharge times (Miao et al., 2023). While a small number of such cases can be normal in the workflow (registrations done on preadmission or discharge decision prior to lab result being returned by the laboratory), a large number of such cases might point to date shifts during extraction. Generally temporal leaks will impact the model performance, with the impact varying depending on how much time has been leaked. The primary way to safeguard from errors is rigorous testing of the ETL process. Manually checking a random sample of the admissions for the sequentially of events (Tomašev et al., 2021) (e.g.: items extracted from ICU system being present before ICU admission) can detect problems not detected during ETL testing. Building a tentative model on the training data and inspecting feature importance can reveal high impact temporal leaks. Timeliness - currency Currency
Temporal leaks due to incorrect date/time handling or incorrect linking during data preparation. Incorrect time zone handling, incorrect joins, time based filtering or grouping and ordering items temporally can introduce temporal leaks during data preparation. Thorough unit testing of the data preparation pipeline can safeguard against temporal leaks bugs. For example, if data are processed for renewing predictions at different event triggers or at equally spaced timepoints (every 6 hours or every 24 hours), testing that processing a full admission will result in the same output as when processing the data linked to the admission up to each trigger time during the admissions. Timeliness - currency Currency
{longtabu}
@@
to >p2cm>p5cm>p5cm>p1cm>p1cm DATA CLEANING Recommendations. EHR: Electronic Health Record, ETL: Extract, Transform, Load, ICU: Intensive Care Unit
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
@@
Non-numeric values in a numeric field Non-numeric values in a numeric field can encode information about the approximate numerical value. Discarding these (or implicitly transforming such fields to numeric and automatically creating missing values) would result in loss of information. A typical example is laboratory values for which machine cap at a specific value, generally because of imprecise measurement capabilities beyond this value: e.g.: D-dimer values of ’> 7650’. The capped value should be extracted as is and preserved as numeric during data preparation, as it contains more information than a missing value. Measurement process - device error - accuracy Correctness
@@
Values outside a physiologically plausible range (outliers) Manually entered numerical values can contain errors. The amount and magnitude of errors depends on the field definition in the EHR software, e.g. some software might not allow values of body temperature higher than 45 °C or dangerous drug doses, displaying an error or warning on the screen when recording such values, while others will allow any numerical value. Exploration of feature ranges and outliers should be done during data preparation. For all numerical values (e.g.: laboratory results, vital signs), a physiologically plausible range can be defined in the data preparation pipeline, and such values can be deleted, or replaced with missing values to be imputed at a later stage in the pipeline. This procedure is called outlier capping (Tomašev et al., 2021). We do not recommend manually correcting such errors, or applying complex rules that cannot be programmatically reproduced at clinical implementation. Measurement process - device error - accuracy; Measurement process - human-introduced error - outliers; Consistency - logical consistency - plausibility Correctness; Plausibility
Clinically implausible combinations of patient characteristics Even when values are within a plausible range for the entire cohort, they can be implausible for the specific patient. Examples include: length of 1m for a neonate; weight of 200g for an adult (Schwabe et al., 2024), patient age at admission is 50 years but admission type ’new-born’ (Miao et al., 2023), paediatric conditions for adults (Blacketer et al., 2021). Such examples will occur mostly in manually entered data. Exploration of feature ranges should be performed also within relevant subgroups for which the plausible ranges differ significantly. As EHR data extraction include a rather large sample size and these errors are rare, these can be set to missing and imputed at a later stage, provided that the imputation model uses relevant features in imputation (e.g.: imputing length using weight and age); mean or median imputation (over all population) might prove as error prone as the initial error. Manual correction of such errors case-by-case would not provide a consistent correction strategy at model implementation time. Correction schemes that can be programmatically reproduced (like imputation) are preferred. Measurement process - human-introduced error - carelessness; Consistency - logical consistency - plausibility Correctness; Plausibility
@@
Implausible values within a time series Numeric measurements within a patient over time constitute time series and outliers can be explored in the context of previous or future values of the time series. For example, slipped sensors used in ICU (Maletzky et al., 2022), as for example a slipped bladder catheter, can record body temperatures closer to room temperature for a period of time, e.g.: 32 °C. While such values are physiologically valid for patients on targeted temperature management, they can be implausible in a specific patient context if they occur only for some minutes or hours. Corrections within a time series should be done during data preparation in a programmatic way, using only previous values from the same time series or from other features. For the temperature use case, a pragmatic approach can be to use only maximum values of temperatures within a time window and use a large enough time window, like 24 hours, within which it is unlikely to have only wrong temperature except if the patient is on targeted temperature management. Measurement process - device error - accuracy; Measurement process - human-introduced error - outliers; Consistency - logical consistency - plausibility Correctness; Plausibility
Inconsistencies between patient admissions Race, sex and age inconsistencies within patient but across several admissions can occur, e.g.: for one admission the race is Hispanic, for another Caucasian (Miao et al., 2023). Presumably, these should be very rare. Depending on the frequency and estimated impact of such inconsistencies on the final model, a decision to correct these in a programmatic way (e.g.: using only past patient admissions data) or leave as is will be context dependent. Measurement process - human-introduced error - carelessness & noisy labels; Consistency - logical consistency - plausibility Correctness; Plausibility
Syntactic variability for missing values Unknown values at the time of registration can be either left as an empty field, or entered as a free text with various values, such as ’Not Mapped’, ’Unknown’, ’Not specified’, ’N/A’, ’NA’, ’?’, etc. (Miao et al., 2023; Ferrao et al., 2016). The values entered in a field should be explored and all values that represent an unknown value at the time of recording should be encoded as missing. Measurement process - human-introduced error - noisy labels; Consistency - rule-based consistency - syntactic consistency Correctness
@@
Character encoding problems Character encoding problems, e.g.: Arteriële Bloeddruk, can either be recorded as such in the EHR fields, due to a bug in the EHR software (that could have manifested only for a limited period of time) or can constitute a bug in the ETL process. Character encoding consistency in the ETL process should be ensured by unit/integration tests. Whenever the problem manifested in the EHR itself, correct mapping to the correctly encoded terminology should be foreseen in the data extraction or preparation. Measurement process - human-introduced error - carelessness; Consistency - rule-based consistency - syntactic consistency Correctness
@@
Inconsistent or incorrectly specified units Units for height, weight, lab results, and other numerical data (Miao et al., 2023; Ferrao et al., 2016) can differ for different measurements, e.g.: height in cm or in m in function of patient’s age, or over time, e.g.: blood glucose in mg/dL and mmol/L. In some EHR systems, units might be specified but might not correctly reflect the units used historically in past years of the data (if the relational database does not store a history of the linked units). Whenever units are correctly specified for all measurements, unit conversion can easily be applied during data preparation. Historical units inconsistencies can be detected by exploration of statistics over time (e.g.: minimum, maximum, mean values) and corrected only for historical data (assuming the units recorded now will be maintained in future data). Measurement process - source credibility - traceability; Consistency - rule-based consistency - syntactic consistency Correctness
@@
Missing values that can be fully recovered from other fields Some extracted data fields might encode, for specific values of the fields, similar or identical information. Admission ward, admission source and admission type are such examples. Missing values in one of these can, in some cases, be recovered from the other extracted fields. Miao et al (Miao et al., 2023) encountered a situation of trivial imputation scenario, in which admission source was emergency room but admission type was missing. In this case, the admission type is also emergency. Whenever the imputation is trivial and easy to reproduce programmatically, this can be performed during data preparation. At the same time, a powerful missing data imputation algorithm might be able to detect and correctly impute such values, given enough correct examples exist in the extracted data. Measurement process - completeness Completeness
@@
Implausible missingness rates While EHR data will inherently exhibit missing values, missingness can be also an artifact of errors during data extraction or preparation. For dynamic prediction models, an exploration of missing values can be done using as a unit of aggregation a sensible choice for communication with clinical experts, such as a patient day (even if the desired aggregation time window is different). This is typically done after feature engineering, and can, in certain situations, reveal feature engineering problems. For example, a missingness rate of around 30% for patient age (which is known in EHR for almost all patients) can have as root cause a problem during linking identifiers, or even in the ETL anonymization process. No missing data in features in which missing data are expected can be due to an unintentional imputation during the ETL process. Missing data can also occur because of incorrect mappings during extraction. For example, if missingness rates in vital signs in ICU, where these are measured at highest frequency, are higher than in other wards, this could prove to be due to incorrect (or omitted) mapping during data extraction or preparation. Whenever the missingness rate is not in line with EHR capabilities or with the expected recording frequency in different wards or patient subgroups, the extraction and preparation logic of those features should be verified for errors. Measurement process - completeness Completeness

### v2

@@ 4.4 Data cleaning
Timestamps when data are available in the system for real-time prediction differ from the clinically relevant timestamp DESCRIPTION: The clinically relevant timestamp reflects the time that best corresponds to the patient’s actual physiological state, and can differ significantly from the data recording timestamp. For laboratory (and microbiology) results (Sauer et al., 2022; Tomašev et al., 2021), the relevant timestamps in the clinical workflow are: order time (when the clinician orders a lab test), planned collection time, collection execution time (when the nurse collects the sample), preliminary result availability time and final (validated) result availability time. Typically, the sample collection time and final result time are extracted from the system. For example, the time when a positive microbiology result is available in the laboratory system is often more than one day after the blood collection by the nurse; the latter is the clinically relevant timestamp because the result reflects the patient state at blood collection time (one day earlier than the result availability). Using the wrong timestamp in prediction will create temporal leaks, but its impact on the individual predictions will differ depending on how features are aggregated in time windows and how often the predictions are renewed (e.g., predictions renewed hourly, every 24 hours or trigger based - when new data become available in the system). Ward transfers can also be entered in the system before or after the actual patient care starts, which represent the clinically relevant time, e.g., transfer to ICU differs from the start of patient monitoring (de Kok et al., 2024).
RECOMMENDATION: Most of the times (and especially for lab results), the relevant timestamps can be extracted from the system. Whenever not possible, the extraction documentation should mention the meaning of the extracted timestamp and its usual recording procedure and possible delays in recording. To avoid time leaks at implementation time, the registration time in the system (result time for lab results) should be used for features included in the prediction model. On the contrary, for outcomes calculated retrospectively based on extracted lab results (e.g., AKI, sepsis, CLABSI), the clinically relevant timestamp (sample collection time for lab results) should be used for outcome calculation, as mentioned in the Table 5.
METRIC: Timeliness - currency; Documentation; WW: Currency
Extracted timestamps may reflect the time of the last modification rather than the original creation time of an item. DESCRIPTION: Extracted timestamps may correspond to the creation time if the item has never been modified or to the last modification time if changes occurred. Most EHR systems do not maintain a full history of modifications, and even when they do, this information is rarely extracted. For example, lab results go through multiple updates: preliminary result, updated preliminary result, validation by a technologist, and final validation by a biologist. While the result value rarely changes, modifications are possible, and updates typically occur within minutes or hours. In contrast, demographic data, such as a patient’s home address, may change over months or years. Since only the most recent address and its last modification timestamp are usually extracted, historical changes might be missing. If spatial exposures are relevant to a prediction task, and extracted data spans over multiple years in which changes likely happened, important information might be missing.
RECOMMENDATION: Extraction documentation should clearly reflect the meaning of extracted timestamps, e.g., creation / modification / result validation time. Insights from system users on how frequently modifications occur can help researchers determine whether the extracted data accurately reflects historical records.
METRIC: Timeliness - currency; Documentation; WW: Currency
@@
No date or timestamps available DESCRIPTION: Some items are not extracted with a timestamp or have date but no time. For example, ICD codes are linked to admission and usually not timestamped; pharmacy medication orders have a date associated, but no timestamp.
RECOMMENDATION: A conservative approach is to encode the timestamp at the end of the relevant time unit (e.g., ICD codes at the end of the admission (Tomašev et al., 2021); medication orders end of day) or a less conservative approach by assigning a typical time of registration (e.g., a very high percentage of medications are ordered before noon, assign 12:00 PM).
METRIC: Timeliness - currency; WW: Currency
Table 9: FEATURE ENGINEERING Recommendations (time-sensitive) Related to Temporal Leaks and Timing Errors. EHR: Electronic Health Record; ETL: Extract, Transform, Load; CPT: Current Procedural Terminology; ICU: Intensive Care Unit; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
The extracted timestamp does not reflect the time when data are recorded in the system. DESCRIPTION: The precise timestamp when items are recorded (or even modified) in the system might not be extracted, or might not be stored in the EHR database for extraction. Typical timestamps differences occur for nurse and clinician observations and coded procedures. For example, nurses might record observations in the patient chart after they finish a shift and they might record an approximation of the time they observed the patient, e.g., at 10:13 the nurse inserts a wound care observation with the time of observation: 8:20 (an estimation of the time the nurse visited the patient). Coded procedures (CPT codes), such as catheter insertion, can be recorded at or after the time of the procedure’s completion and the timestamp assigned to the procedure is the execution time or a close approximation of it. However, CPT codes can also be reviewed and documented retrospectively, even after patient discharge, when coding is reviewed for billing, based on patient’s documentation throughout the hospitalization. There can be variation in the recording process, even within the same procedure code: sometimes procedures codes are recorded at the time of execution and other times after patient discharge. The discrepancy between the data available at model development and data available at prediction time in the clinical implementation may be larger for coded procedures compared to the nurse tasks, because of their review for billing purposes.
RECOMMENDATION: The extraction documentation should mention the meaning of the extracted timestamp and its usual recording procedure and possible delays in recording. These are though difficult to document exhaustively, as the nurse flow might differ in different wards and coding procedures might vary widely, situations in which understanding of the data recording procedures will require communication with the clinical experts.
METRIC: Timeliness - currency; Documentation; WW: Currency
Exact-hour timestamps often do not reflect the time when data are recorded in the system DESCRIPTION: If a large proportion of timestamps in a date/time field appear as exact hours (e.g., 08:00:00 or 23:00:00), it may indicate that the extracted timestamps deviate from the time of recording in the system. Nurses might complete the tasks while they are at patient’s bedside (09:12:30), but they record them in the EHR software after they finish their round (at 11:31:30), defaulting the execution time to a pre-set planned time, such as the start of a nurse’s shift or a standard scheduled time (08:00:00), instead of the time the task was executed in the system. In this example, the extracted timestamp might be 08:00:00, while the time available in the system for prediction would be 11:31:30. Depending on the EHR software and the hospital flow, an exact-hour timestamp could also mean that the task was planned, but not executed (when task execution status is not available in the extraction).
RECOMMENDATION: The data extraction documentation of the extracted timestamp columns should mention the exact meaning of the extracted timestamps: execution, planning or registration time, or a combination thereof. Although the procedures for data recording might differ from ward to ward, whenever possible, the hospital flow should be documented and understood. When the exact-hour timestamps do not represent the meaningful timestamps, but the clinical items recorded are of predictive value, encoding the timestamps at the end of the relevant time unit (e.g., end of nurse shift) for model training would help avoiding temporal leaks.
METRIC: Timeliness - currency; Documentation; WW: Currency
@@
Temporal leaks due to extraction errors DESCRIPTION: Temporal leaks can also occur because of errors during the ETL process, for example introduced by incorrect time zone handling, incorrect joins or time based filtering clauses. These are typically difficult to spot during data preparation. If the temporal leaks have a profound effect on the prediction performance, features that are not expected to have very high importance might appear as most important (e.g., history of bacteremia being the most important feature in predicting bacteremia during current admission might point to incorrectly leaking the current admission in the extraction of the historical feature). Some measurement timestamps may fall outside the admission period, occurring either before admission or after discharge. (Miao et al., 2023). While a small number of such cases can be normal in the workflow (registrations done on preadmission or discharge decision prior to lab result being returned by the laboratory), a large number of such cases might point to date shifts during extraction. Generally temporal leaks will impact the model performance, with the impact varying depending on how much time has been leaked.
RECOMMENDATION: The primary way to safeguard from errors is rigorous testing of the ETL process. Manually checking a random sample of the admissions for the sequentially of events (Tomašev et al., 2021) (e.g., items extracted from ICU system being present before ICU admission) can detect problems not detected during ETL testing. Building a tentative model on the training data and inspecting feature importance can reveal high impact temporal leaks.
METRIC: Timeliness - currency; WW: Currency
Temporal leaks due to incorrect date/time handling or incorrect linking during data preparation. DESCRIPTION: Incorrect time zone handling, incorrect joins, time based filtering or grouping and ordering items temporally can introduce temporal leaks during data preparation. Such leaks can distort predictions by providing information that would not be available at the time of prediction in clinical implementation. For example, misalignment of time zones may shift timestamps incorrectly, making future events appear as if they occurred earlier. Incorrect joins between tables (e.g., linking lab results to patient records without proper date/time filtering) can introduce data that should not be available at prediction time.
RECOMMENDATION: Thorough unit testing of the data preparation pipeline is essential to prevent temporal leaks. For example, when predictions are generated at specific triggers (every 6 or 24 hours), testing should ensure that processing a full admission retrospectively in the data pipeline produces the same results as processing data incrementally up to each trigger time. This step helps confirm that only information available at a given prediction time is used, preserving the model’s real-world applicability.
METRIC: Timeliness - currency; WW: Currency
@@
Table 10: DATA CLEANING Recommendations Related to Outliers and Encoding Issues. EHR: Electronic Health Record, ETL: Extract, Transform, Load; ICU: Intensive Care Unit; CPR: cardiopulmonary resuscitation; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
@@
Non-numeric values in a numeric field DESCRIPTION: Non-numeric values in a numeric field can encode information about the approximate numerical value. Discarding these (or implicitly transforming such fields to numeric and automatically creating missing values) would result in loss of information. A typical example is laboratory values that are capped at a specific threshold due to the detection limit of the assay, meaning measurements beyond this value are not precisely quantified (Albert, 2008). For instance, D-dimer results may be reported as ’> 7650’ when they exceed the assay’s measurable range.
RECOMMENDATION: The capped value should be extracted as is and preserved as numeric during data preparation, as it contains more information than a missing value.
METRIC: Measurement process - device error - accuracy; WW: Correctness
@@
Values outside a physiologically plausible range (outliers) DESCRIPTION: Manually entered numerical values can contain errors. The amount and magnitude of errors depends on the field definition in the EHR software, e.g. some software might not allow values of body temperature higher than 45 °C or dangerous drug doses, displaying an error or warning on the screen when recording such values, while others will allow any numerical value.
RECOMMENDATION: Exploration of feature ranges and outliers should be done during data preparation. For all numerical values (e.g., laboratory results, vital signs), a physiologically plausible range can be defined in the data preparation pipeline, and such values can be deleted, or replaced with missing values to be imputed at a later stage in the pipeline. This procedure is called outlier capping (Tomašev et al., 2021). We do not recommend manually correcting such errors, or applying complex rules that cannot be programmatically reproduced at clinical implementation.
METRIC: Measurement process - device error - accuracy; Measurement process - human-introduced error - outliers; Consistency - logical consistency - plausibility; WW: Correctness; Plausibility
@@
Implausible or inconsistent values within a time series DESCRIPTION: Numeric measurements within a patient over time constitute time series and outliers can be explored in the context of previous or future values of the time series. For example, slipped sensors used in ICU (Maletzky et al., 2022), as for example a slipped bladder catheter, can record body temperatures closer to room temperature for a period of time, e.g., 32 °C. While such values are physiologically valid for patients on targeted temperature management, they can be implausible in a specific patient context if they occur only for some minutes or hours. Another example constitutes a heart rate of 0 used to mark admissions not linked to a real patient (de Kok et al., 2024), while in the ICU a heart rate of 0, within the context of previous normal heart rate followed by a normal heart rate can indicate a CPR procedure was done. Race, sex and age inconsistencies within patient but across several admissions can occur, e.g., for one admission the race is Hispanic, for another Caucasian (Miao et al., 2023). Presumably, these should be very rare.
RECOMMENDATION: Corrections within a time series should be done during data preparation in a programmatic way (automated, reproducible, re-executable at prediction time), using only previous values from the same time series or from other features. For the temperature use case, a pragmatic approach can be to use only maximum values of temperatures within a time window and use a large enough time window, like 24 hours, within which it is unlikely to have only low temperature values except if the patient is on targeted temperature management. If inconsistencies are rare and a programmatic way of correcting these proves unfeasible (e.g., is the patient truly Caucasian or truly Hispanic?), a decision leave as is might be the best option.
METRIC: Measurement process - device error - accuracy; Measurement process - human-introduced error - outliers; Consistency - logical consistency - plausibility; WW: Correctness; Plausibility
Character encoding problems DESCRIPTION: Character encoding problems, e.g., Menière’s disease, can either be recorded as such in the EHR fields, due to a bug in the EHR software (that could have manifested only for a limited period of time) or can constitute a bug in the ETL process.
@@
RECOMMENDATION: Character encoding consistency in the ETL process should be ensured by unit/integration tests. Whenever the problem manifested in the EHR itself, correct mapping to the correctly encoded terminology should be foreseen in the data extraction or preparation.
METRIC: Measurement process - human-introduced error - carelessness; Consistency - rule-based consistency - syntactic consistency; WW: Correctness
@@
Inconsistent or incorrectly specified units DESCRIPTION: Units for height, weight, lab results, and other numerical data (Miao et al., 2023; Ferrao et al., 2016) can differ for different measurements, e.g., height in cm or in m in function of patient’s age, or over time, e.g., blood glucose in mg/dL and mmol/L. In some EHR systems, units might be specified but might not correctly reflect the units used historically in past years of the data (if the relational database does not store a history of the linked units).
RECOMMENDATION: Whenever units are correctly specified for all measurements, unit conversion can easily be applied during data preparation. Historical units inconsistencies can be detected by exploration of statistics over time (e.g., minimum, maximum, mean values) and corrected only for historical data (assuming the units recorded now will be maintained in future data, at implementation).
METRIC: Measurement process - source credibility - traceability; Consistency - rule-based consistency - syntactic consistency; WW: Correctness
Table 11: DATA CLEANING Recommendations Related to Missing Data and Logical Inconsistencies. EHR: Electronic Health Record; ETL: Extract, Transform, Load; ICU: Intensive Care Unit; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
Clinically implausible combinations of patient characteristics DESCRIPTION: Even when values are within a plausible range for the entire cohort, they can be implausible for the specific patient. Examples include: length of 1m for a neonate; weight of 200g for an adult (Schwabe et al., 2024), patient age at admission is 50 years but admission type ’new-born’ (Miao et al., 2023), paediatric conditions for adults (Blacketer et al., 2021). Such examples will occur mostly in manually entered data.
RECOMMENDATION: Exploration of feature ranges should be performed also within relevant subgroups for which the plausible ranges differ significantly. As EHR data extraction include a rather large sample size and these errors are rare, these can be set to missing and imputed at a later stage, provided that the imputation model uses relevant features in imputation (e.g., imputing length using weight and age); mean or median imputation (over all population) might prove as error prone as the initial error. Manual correction of such errors case-by-case would not provide a consistent correction strategy at model implementation time. Correction schemes that can be programmatically reproduced (like imputation) are preferred.
METRIC: Measurement process - human-introduced error - carelessness; Consistency - logical consistency - plausibility; WW: Correctness; Plausibility
Syntactic variability for missing values DESCRIPTION: Unknown values at the time of registration can be either left as an empty field, or entered as a free text with various values, such as ’Not Mapped’, ’Unknown’, ’Not specified’, ’N/A’, ’NA’, ’?’, etc. (Miao et al., 2023; Ferrao et al., 2016).
RECOMMENDATION: The values entered in a field should be explored and all values that represent an unknown value at the time of recording should be encoded as missing.
METRIC: Measurement process - human-introduced error - noisy labels; Consistency - rule-based consistency - syntactic consistency; WW: Correctness
@@
Missing values that can be fully recovered from other fields DESCRIPTION: Some extracted data fields might encode, for specific values of the fields, similar or identical information. Admission ward, admission source and admission type are such examples. Missing values in one of these can, in some cases, be recovered from the other extracted fields. Miao et al (Miao et al., 2023) encountered a situation of trivial imputation scenario, in which admission source was emergency room but admission type was missing. In this case, the admission type is also emergency.
RECOMMENDATION: Whenever the imputation is trivial and easy to reproduce programmatically, this can be performed during data preparation. At the same time, a powerful missing data imputation algorithm, that effectively utilizes patterns in other features, might be able to detect and correctly impute such values, given enough correct examples exist in the extracted data.
METRIC: Measurement process - completeness; WW: Completeness
@@
Implausible missingness rates DESCRIPTION: While EHR data will inherently exhibit missing values, missingness can be also an artifact of errors during data extraction or preparation. For dynamic prediction models, missing data patterns can be explored using an aggregation unit that is meaningful for communication with clinical experts, such as a patient day, even if the final model uses a different aggregation time window. This is typically done after feature engineering, and can, in certain situations, reveal feature engineering problems. For example, a missingness rate of around 30% for patient age (which is known in EHR for almost all patients) can have as root cause a problem during linking identifiers, or even in the ETL anonymization process. No missing data in features in which missing data are expected can be due to an unintentional imputation during the ETL process. Missing data can also occur because of incorrect mappings during extraction. For example, if missingness rates in vital signs in ICU, where these are measured at highest frequency, are higher than in other wards, this could prove to be due to incorrect (or omitted) mapping during data extraction or preparation.
RECOMMENDATION: Whenever the missingness rate is not in line with EHR capabilities or with the expected recording frequency in different wards or patient subgroups, the extraction and preparation logic of those features should be verified for errors.
METRIC: Measurement process - completeness; WW: Completeness